There were 610 people at a play. The admission price was $3.00 for adults and $1.00 for children. The admission receipts were $1250. How many adults and children attended?
I got the answer of 290 adults/320 children. Is this correct?
Thanks!

Answer by nerdybill(7384)

There were 610 people at a play. The admission price was $3.00 for adults and $1.00 for children. The admission receipts were $1250. How many adults and children attended?
.
You have it reversed...
.
Let x = number of adult tickets
then
610-x = number of children tickets
.
3x + 1(610-x) = 1250
3x + 610-x = 1250
2x + 610 = 1250
2x = 640
x = 320 (adults)
.
children:
610-x = 610-320 = 290
.